LOANS: Written Answer Questions (3) Due 7/2 1 Sylvia is repaying a loan of X by making 17 annual payments of 100. Each paymenall of three please
Answer #1

Loan amount=X

Effective interest rate=5%

Number of payments=17

Annual Payment=100

Annual interest at 5%=X*0.05

Principal deposited in sinking fund each year=100-0.05X

Compound Amount Factor(CAF)=(F/A,i,N)=(((1+i)^N)-1)/i

i= Interest rate in sinking fund =4%=0.04

N=Number of Years=17

Compound Amount Factor(CAF)=(F/A,4%,17)=(((1+0.04)^17)-1)/0.04=23.69751

Loan Balance at end of 17 Years=X

Sinking fund accumulated at sinking fund rate=23.69751*(100-0.05X)

23.69751*(100-0.05X)=X

2369.751-1.184876X=X

X*(1+1.184876)=2369.751

X=2369.751/2.184876=1084.616

X=1085 (Rounded to whole number)